MC 100! Demand 65 100 200 Number freeway users

PART A)

Consider Amy, who has an easy access to the metro, meaning that her best alternate route is metro. Assuming that Amy’s cost using the metro is $60, she is a

a) freeway user in equilibrium but a metro user in social optimum.

b) metro user both in equilibrium and in social optimum.

c) freeway user both in equilibrium and in social optimum.

d) metro user in equilibrium but a freeway user in social optimum.

PART B)

Which of the followings is the socially optimal congestion toll imposed per person under the equilibrium with tolling?

a)32

b)10

c)15

d)50

e)100

f) 65

PART C)

Calculate the toll revenue.

2080

a)1500

b)975

c)3200

d)1105

PART D)

Calculate the welfare loss (deadweight loss) that arises due to externalities from congestion.

a)875

b)4225

c)2112.5

d)1040

e)5000
Answer #1

A.

Correct Answer:

D

The cost of ride in ,metro is less than the cost of use of the freeway at a social equilibrium. So, Amy is in equilibrium with metro, but in socially optimal equilibrium with freeway.

======

B.

Correct Answer:

C

The size of the toll should be equal to the marginal external cost (MEC).

MEC = 65-50

MEC = $15

======

C.

Correct Answer:

B

Toll revenue = (65-50)*65

Toll revenue = $975

=======

D.

Correct Answer:

A

Dead weight loss = 0.5*(100-65)*(100-65) + 0.5*(100-65)*(65-50)

Dead weight loss = $875